OpenStudy (anonymous):

simplify x/6x-x^2 simplify -12x^4/x^4+8x^5

OpenStudy (***[isuru]***):

\[\frac{ x }{ 6x - x^{2} }\]\[\frac{ x }{ x( 6 - x) }\]\[\frac{ 1 }{ 6- x }\]

OpenStudy (***[isuru]***):

\[\frac{ -12x^{4} }{ x^{4} + 8x^{5} }\]\[\frac{ -12x^{4} }{ x^{4}(1 - 8x) }\]\[\frac{ -12 }{ (1-8x) }\]\[\frac{ 12 }{ (8x - 1) }\]

OpenStudy (anonymous):

where x = 0,6 or where x=6?

OpenStudy (***[isuru]***):

nope... x can't be equal to 6 'cause then the expression become impossible to define as the denominator becomes 6... so i think it's where x is not equal to 6

OpenStudy (anonymous):

ohhhh kay thank youuuu ( :

OpenStudy (***[isuru]***):

u r welcome ! ;)